OpenStudy (anonymous):

A line has a rise of 6 and a slope of 1/20. What is the run?

OpenStudy (anonymous):

slope = rise/run 6/x=1/20 x=120 The run is 120 :P

OpenStudy (saifoo.khan):

Solve for run.